A plate 0.025 mm distant from a fixed plate, moves at 50 cm/s and requires a force of 1.471 N/m^2 to maintain this speed. Determine the fluid viscosity between the plates in the poise.
1
Expert's answer
2020-11-11T20:15:20-0500

Distance between plates


"dy=0.025mm=2.5\times10^{-5}m"

Velocity of upper plate


"u=50cm\/s=0.5m\/s"

Force on upper plate


"F=1.471\ N/m^2=\tau"

Hence

The value of the share stress


"\tau=1.471\ N/m^2"

Change of velocity


"du=u-0=0.5m\/s"

Change of distance


"dy=2.5\times10^{-4}m"

"\tau=\mu\cdot\dfrac{du}{dy}"

where "\mu" is the fluid viscosity between the plates


"\mu=\dfrac{\tau\cdot dy}{du}"

"\mu=\dfrac{1.471\ N/m^2 \cdot 2.5\times10^{-5}m}{0.5m/s}="

"=7.355\times10^{-5}N\cdot s/m^2=7.355\times10^{-4} \text{poise}"

"7.355\times10^{-4}" poise